OpenStudy (raden):

Find all triples (x,m,n) of positive integers satisfying the equation x^m=2^{2n+1}+2^n+1.

OpenStudy (anonymous):

\[\large{x^m=2^{2n+1}+2^n+1}~?\]

OpenStudy (raden):

yes

OpenStudy (anonymous):

Let \(u=2^n\). Then you have a quadratic equation in terms of u: \(2u^2+u+(1-x^m)=0\). Hope that helps.

OpenStudy (raden):

next, what should i do??

OpenStudy (anonymous):

I really have no idea. But \(\large 1-x^m=(1-x)(x^{m-1}+x^{m-2}y+\ldots +xy^{m-2}+y^{m-1})\). Sooo... Since u is rational, perhaps apply the rational root theorem?

OpenStudy (raden):

hmmm, thanks for ur clue... but i think very" difficul for me i hope anyone give me more to solve this problem.. thanks @Herp_Derp

OpenStudy (anonymous):

dont worry i'll solve it soon ;)

OpenStudy (anonymous):

lol ; not that easy...solved it before for m=2 but for m>2 ........

OpenStudy (experimentx):

running a code for half and hour ... still no more than two solutions n=0, n=4

OpenStudy (anonymous):

I believe that only answers are \( (x,m,n)=(2^{2k+1}+2^k+1,1,k) \ and \ (23,2,4) \) but my proof is still incomplete......

OpenStudy (experimentx):

for n>0 it would be that for n=0 (2, 2, 0) this is also true. also note that m>1

OpenStudy (anonymous):

m,n and k are >0 positive integers

OpenStudy (experimentx):

ah ... it means i didn't read the question well

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i have the complete proof now but it takes me an hour to type it....

OpenStudy (anonymous):

Well note that \(x\) is an odd number. \(m=1 \) is an obvious answer so i want to work on \(m>1 \).............. firstly let \(m=2k+1 \) where \(k \ge 1 \) we have \[ 2^{n}(2^{n+1}+1)=x^{m}-1=(x-1)(x^{m-1}+...+1)\]\( (x^{m-1}+x^{m-2}+...+x+1)=(x^{2k}+x^{2k-1}...+x+1) \) is an odd number because \(x\) is odd so \(x-1\) divides \(2^n\) then \(x-1 \ge 2^n\) and \(x\ge 1+2^n\) hence : \[ 2^{2n+1}+2^{n}+1=x^{m}\geq (2^{n}+1)^{3}=2^{3n}+3\times 2^{2n}+3\times 2^{n}+1 \]and this is a contradiction..... now let \(m=2k\) then \( 2^{n}(2^{n+1}+1)=x^{2k}-1=(x^k-1)(x^k+1)=(y-1)(y+1) \) and \(y\) is odd. \(y-1\) and \(y+1\) are even factors so \( \gcd(y-1,y+1)=2 \) exactly one of them divisible by \(4\). Hence \(n ≥ 3\) and one of these factors is divisible by \(2^{n−1}\) but not by \(2^{n} \). So we can write \(y = 2^{n−1}s +t\) , \(s\) odd , \(t= ±1\). Plugging this into the original equation we obtain \(2^n(2^{n+1}+1)=(2^{n−1}s +t)^2-1=2^{2n−2}s^2+2^n st \) or, equivalently \(2^{n+1}+1=2^{n−2}s^2+ st \) Therefore \(1-st=2^{n−2}(s^2-8) \) For \(t= 1\) this yields \(s^2 − 8 ≤ 0\), so \(s = 1\), which fails to satisfy original equation. For \(t=-1\) equation gives \(1+s=2^{n−2}(s^2-8) \ge 2(s^2-8)\) so \(2s^2-s-17 \le 0\) and \( s \le 3\) we know that \(s\) is odd so \(s=3\). put this in the last equation with \(t=-1 \) gives \(n=4\). back to the original equation with \(n=4\) u have \[x^m=529=23^2 \] so \(x=23\) and \(m=2\) the only solutions are \( (x,m,n)=(2^{2l+1}+2^l+1, 1,l) \ and \ (23,2,4) \)

OpenStudy (experimentx):

man that's a long solution!!

OpenStudy (anonymous):

yes... i've learned it for m=2 from a book or something; i cant remember... it helped me a lot

OpenStudy (experimentx):

great!!

OpenStudy (anonymous):

a little typo *****which fails to satisfy *last* equation.******

OpenStudy (anonymous):

@RadEn come here plz

OpenStudy (raden):

hoho, yea... wht about u with other solution in inbox (emai)

OpenStudy (anonymous):

but those are incorrect...check them... (11,1,1), (37, 1, 2) are incorrect

OpenStudy (anonymous):

im pretty sure that there is no other solution for this diophantine equation

OpenStudy (raden):

but, for (x,m,n)=(11,1,1) => 11^1= 2^(2(1)+1) + 2^1 + 1 11 = 8 + 2 + 1 i think Right side = Left side

OpenStudy (anonymous):

oh so thats answers for m=1 i provided a closed form for particular case of m=1

OpenStudy (raden):

my friend tell me that the solution has infinitely, but i cant prove that

OpenStudy (anonymous):

i already proved it for u ...m=1 is an obvious solution of this equation...

OpenStudy (anonymous):

let \(m=1\) u will have \(x=2^{2n+1}+2^n+1 \) now for any natural number like k the answer is \( (x,m,n)=(2^{2k+1}+2^k+1,1,k) \)

OpenStudy (anonymous):

that equation becomes to an obvious form with letting m=1

OpenStudy (raden):

but, if we look the problem original : x^m = 2^(2n+1) + 2^n + 1 satisfy for solution (11,1,1), (37,1,2) isnt it? and the problem tell that all triples (x,m,n) are positive integers

OpenStudy (anonymous):

yes thats right and closed form of this solutions is what i write up there...

OpenStudy (anonymous):

just let k=1,2,3,...

OpenStudy (raden):

so, what conklusy the solutions it,, only one or more??

OpenStudy (anonymous):

there are infinite solution

OpenStudy (anonymous):

solutions are \((x,m,n)=(23,2,4)\) and \((x,m,n)=(2^{2k+1}+2^k+1,1,k)\) \(k=1,2,3,...\)

OpenStudy (raden):

ok, thank u very much.. i want to save this page, n i will study it ... case closed.. hehe

OpenStudy (anonymous):

OK...save the address of page...

OpenStudy (raden):

ok.. see u tomorrow :D